OpenStudy (anonymous):

Given f(x) = the quantity of 3x minus 1, divided by 2, solve for f−1(4). 1 3 5 7

OpenStudy (anonymous):

medal to who ever answers this correctly

OpenStudy (anonymous):

@ganeshie8 @nincompoop

OpenStudy (anonymous):

@muscrat123

OpenStudy (muscrat123):

sowwy, busy w/ flvs. got 50 2 do in 2 days

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (lynfran):

f(x)=(3x-1)/2 y=(3x-1)/2 x=(3y-1)/2 solve for y we get 2x=3y-1 2x+1=3y (2x+1)/3=y f^-1(x)=(2x+1)/3 f^-1(4)=(2(4)+1)/3 ...can u solve it...

OpenStudy (anonymous):

so c is correct?

OpenStudy (lynfran):

no... (2(4)+1)/3 (8+1)/3 9/3=?

OpenStudy (anonymous):

so b @LynFran

OpenStudy (lynfran):

yes @jammy987